Calculate the new force between two masses, if one of the mass is 1/3rd ,that of the other.
1
Expert's answer
2019-09-30T09:49:03-0400

The force will be:


"F=G\frac{m(\frac{1}{3}m)}{r^2}=G\frac{m^2}{3r^2}"